OpenStudy (anonymous):

8. Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?

OpenStudy (anonymous):

let original cost price = x profit = 35% of x = 0.35x selling price = cost price + profit = x + 0.35x = 1.35x ATQ, 1.35x = 670 ==> x = 670/1.35 so x = 496.30

OpenStudy (anonymous):

if u r happy with the ans, pls click GOOD ANSWER button for me o( :{)